Calculate the magnitude of the gravitational force exerted on a cantaloupe of mass m=1 kg on the surface of the earth due to (a)the earth(b)the moon(c) the sun. The distance of the earth from the moon to 3.82×10^8 and the sun is 1.5×10^11.(mass of moon=7.66×10^22 & mass of sum=1.99×10^30)

Grade:12th pass

2 Answers

Venkat
5 years ago
1) m = 1 kg;  M = 5.98 x 1024 kg;  rearth= 6.37 x 106 m;  G = 6.67 x 10-11 m3 / (kg s2)

Fg = GmM/r2
Fg = [(6.67 x 10-11 m3 / (kg s2)) (1 kg)(5.98 x 1024 kg)]/( 6.37 x 106 m)2
Fg = 9.821 N

note this is approximately equal to F = mg = 1 kg(9.81 m/s2) = 9.81 N

2)
Fm = GmM/r2
Fm = [(6.67 x 10-11 ) (1)(7.66 x 1022 )]/(3.82 x 108 )2
Fm = 3.51 x 10-5 N
 

3)
 
Fs = GmM/r2
Fs = [(6.67 x 10-11 ) (1)(1.99 x 1030 )]/(1.5 x 1011 )2
Fs = 0.00589 N
SADDAM HUSSAIN
2 years ago
calculate the magnitude of the gravitational force exerted on a body of mass 1.00kg on the surface of earth due to earth moon sun